OpenStudy (anonymous):

update: the correct question is 7x + 4-9x=12x-48

OpenStudy (anonymous):

x=26/7

OpenStudy (stormfire1):

7x + 4 - 9x = 12x - 48 7x - 12x -9x = -48 - 4 14x = -52 x = -52/14 = -26/7

OpenStudy (anonymous):

thanks!=)